Doc. Math. J. DMV 97

The Work of the Fields Medalists and of the Rolf Nevanlinna Prize Winner

Peter Goddard: The Work of Richard Ewen Borcherds . . . . . . . . . . . . I  99
Béla Bollobás: The Work of William Timothy Gowers . . . . . . . . . . . . I  109
Clifford Henry Taubes: The Work of Maxim Kontsevich . . . . . . . . . I  119
Steve Smale: The Work of Curtis T. McMullen . . . . . . . . . . . . . . . . . . . . I  127
Ronald Graham: The Work of Peter W. Shor . . . . . . . . . . . . . . . . . . . . . . I  133

Documenta Mathematica · Extra Volume ICM 1998 · I

98

Documenta Mathematica · Extra Volume ICM 1998 · I